OpenStudy (anonymous):

Please help!!!!!!!! solve using the quadratic formula.. x^2-7x-6=0

OpenStudy (anonymous):

in this problem what is a , b and c?

OpenStudy (anonymous):

i doesnt say?

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

how do you know whats what?

OpenStudy (anonymous):

When using the quadratic formula \[ax^2+bx+c=0\] is the form. So based on your eqaution \[x^2-7x-6\] what is in the place of a, b and c.

OpenStudy (anonymous):

a=x^2 b=-7 c=6 ?

OpenStudy (anonymous):

close a is the coefficient on x^2 not the x^2 itself. (Do you know what a coefficient is?)

OpenStudy (radar):

(a) is not the variable itself, but the coefficient, which is the numerical value that is being used to multiply the variable, if none is there then a=1.

OpenStudy (anonymous):

Thanks Radar....but the question was for lashell123... :-)

OpenStudy (radar):

Yes, and I want lashell123 to determine "a"

OpenStudy (radar):

It was obvious the post, lashell123 did not understand "1" as a coefficient when it was not shown.

OpenStudy (anonymous):

So do I ...but she seems to be away at the moment.

OpenStudy (anonymous):

And that is why I was walking her through it. Please don't take offense.

OpenStudy (radar):

She has closed the question, but I thought she might still be viewing it, or still interested in solving it.

OpenStudy (radar):

No offense taken, and you were indeed helping her learn.

OpenStudy (anonymous):

:-)